# ChartBlender: An Interactive System for Authoring and Synchronizing Visualization Charts in Video

### v1

@@ Abstract
Embedding data visualizations in video can enhance the communication of complex information. However, this process is often labor-intensive, requiring designers to adjust visualizations frame by frame manually. In this work, we present ChartBlender, a novel system that streamlines this process by enabling users to create data visualizations, embed them seamlessly into video scenes, and automatically synchronize them with both camera motion and moving objects. Particularly, ChartBlender incorporates a tracking algorithm that supports both object and camera tracking, ensuring robust alignment of visualizations with dynamic video content. To maintain visual clarity and aesthetic coherence, we also explore the design space of video-suited visualizations and develop a library of customizable templates optimized for video embedding. We evaluate ChartBlender through two controlled experiments and expert interviews with five domain experts. Results show that our system enables accurate synchronization and accelerates the production of data-driven videos.
Index Terms:
data visualization, video editing, object tracking, camera tracking, interface, visual storytelling.
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ 1 Introduction
Embedding data visualizations into videos has emerged as a powerful method for conveying complex information [1, 2, 3, 4, 5]. Such a form of representation is increasingly used across various domains (e.g., sports, traffic, and journalism), allowing audiences to access supplementary information without being distracted from the primary video content[6]. Despite its potential, the process remains labor-intensive: achieving spatial and temporal synchronization between visualizations and video typically requires manual, frame-by-frame alignment. Even a 10-second visualization-embedded video may demand up to a full day of manual work from designers.
Motivated by a collaborative project with a major Chinese media organization to create a chart-embedded video, we conducted a comprehensive survey of existing tools and systems. On one end of the spectrum are professional tools like Adobe After Effects [7] and Blender[8]. While powerful, these tools rely on manual parameter tuning and lack intelligent assistance, making them difficult to use efficiently. On the other end, recent advances in Augmented Reality (AR) and Virtual Reality (VR) have inspired research systems that allow casual users to overlay visualizations on videos for sharing quickly [9]. These systems prioritize ease of use and rapid output but are designed primarily for end users with simple needs. As a result, they fall short in supporting high-quality, professionally produced content. Specifically, they lack robust motion tracking for objects with diverse movement patterns, as well as precise control over visualization placement and editing capabilities essential for content creators and news organizations.
@@
Therefore, there is a pressing need for an interactive authoring system supports flexible, efficient embedding of data visualizations into complex video scenes and ensures accurate motion synchronization throughout the video. However, to design such an effective system, three challenges need to be tackled: (1) track the motion of videos accounting for both moving objects in the scene and camera pose variations, to ensure consistent synchronization between visualizations and the video scene; (2) design generalized visualizations to support clear and informative presentation of data within dynamic video contexts; and (3) enable user-driven authoring and refinement of visualizations through simple and integrated interaction. Addressing these challenges is essential for the precise and seamless embedding of data visualizations into real-world videos.
@@
To address these challenges, we introduce ChartBlender, a novel authoring system for embedding and automatically synchronizing data visualizations with video during playback. In particular, the user simply selects a target tracking point and a preferred tracking mode, and the system automatically generates the complete motion trajectory for visualization throughout the Users can edit the visualization at any time, including both its visual style and its 3D pose within the video scene. In this way, the system supports the human-machine collaboration to generate data-driven videos rapidly while ensuring the visualization appears visually coherent and seamlessly embedded into the video. The major contributions of the paper are as follows:
@@
  • System. We introduce ChartBlender, a novel authoring system that enables the authoring and automatic synchronization of data visualizations with video content. It features an interactive interface that supports the end-to-end creation of visualization-embedded videos through an integrated authoring workflow.
@@
  • Algorithms. We develop two algorithms to support robust visualization embedding in dynamic video scenes. The first estimates frame-by-frame camera poses, enabling visualizations to remain stably anchored to the background. The second computes and refines object trajectories across frames, allowing visualizations to accurately follow moving targets within the video.
@@ 2 Related Work
In this section, we survey the papers that are most relevant to our work, including professional video editing software, embedded visualization, and point tracking.
@@ 2.1 Video Editing Software
Video editing is a broad field that involves various techniques, technologies, and tools used to create, manipulate, and refine video content, which often pose barriers to users [10, 11]. Professional video editing software offers powerful tools to insert animated objects into videos. Traditionally, users must manually keyframe an object’s position, rotation, and scale over time, with smooth transitions achieved through interpolation. While recent advances have introduced automated tracking methods, such as 3D camera estimation and object tracking, these techniques, available in tools like Adobe After Effects, Blender, and Nuke [7, 8, 12], still require a level of expertise that can be challenging for beginners. At the same time, the growing demand for data-driven videos highlights a gap in current workflows: mainstream video editing tools lack integrated support for data visualization. Users often switch between separate tools or project files to create visualizations, making it difficult to assess whether a chart fits naturally within a given video context. Currently, no tool allows for direct editing of chart styles within the video environment while supporting real-time assessment, motion synchronization, and object tracking. This creates a pressing need for authoring tools that bridge the gap between visualization and video editing.
2.2 Embedded Visualization in Video
Embedded data visualization in videos has become increasingly popular [13, 14, 15, 16]. However, this process often requires users to manually create keyframes for positioning the visualizations within the video. Recently, there has been increasing research on intelligent systems that can help embed visualizations in videos [17, 18, 6, 19]. Early works primarily focused on embedding static charts into videos. For instance, SmartShot [17] developed a system that optimizes the placement of charts to avoid obstructing the visual focus, ensuring aesthetic balance and readability. SwimFlow [18] explored how to embed visualizations into swimming competition videos. Their system allows charts to animate by using manual annotations to mark athletes’ starting positions and interpolating to calculate their motion trajectories. However, this approach is clearly limited when dealing with more complex motion patterns. Recent advancements in computer vision now enable the embedding of visualizations directly into sports videos, offering a more dynamic and adaptive solution. For instance, iBall [6] embeds dynamic visual data cues into basketball videos and automatically adjusts the information display based on the user’s gaze. However, when the camera moves, the charts can only approximate the motion rather than being stably anchored to the scene. Additionally, these methods typically involve 2D overlays and tracking.
In contrast to these domain-specific solutions, our approach tracks 3D trajectories of arbitrary objects and supports 3D effects, including perspective transformations and occlusions of the inserted visualizations. Moreover, existing systems often lack the flexibility to freely create and edit visualizations to meet user-specific needs. In our system, users can directly edit the types, styles, and data mappings of charts on the video, streamlining the creation process and offering greater customization.
@@ 2.3 Point Tracking
Tracking arbitrary points in videos has witnessed substantial progress in recent years [20, 21, 22], with methods such as TAP-Net [23] demonstrating the ability to reliably follow both static and dynamic points throughout a video sequence. These approaches typically leverage appearance-based cues and temporal consistency to maintain accurate correspondences, even under significant motion and occlusion. However, focusing solely on 2D tracking is often insufficient in practical applications. While a dynamic point in the scene can be tracked frame-by-frame, certain use cases, such as visual overlays or interactive annotations, require this point to remain anchored relative to the static scene background, rather than to moving objects or transient visual cues. To achieve this, it becomes necessary to disentangle camera motion from scene dynamics, which calls for robust camera pose estimation.
Fortunately, a variety of mature techniques now exist for estimating camera pose from monocular or multi-view video streams [24, 25], ranging from classical SLAM-based pipelines [26, 27, 28] to deep learning methods[29, 30, 31, 32, 33]. The integration of 2D point tracking with camera pose estimation enables a more unified paradigm: tracking arbitrary points in 3D space. This formulation has recently attracted increasing attention [34, 35, 36], as researchers extend traditional 2D tracking frameworks into the 3D domain. For instance, SpatialTracker[36] is the first work specifically designed for the task of 3D point tracking. It estimates the 3D trajectory of arbitrary pixels by combining image features with monocular depth estimation, and propagates information through a graph neural network (GNN) to model long-range spatiotemporal motion. Similarly, TAPIR-3D[34] is an experimental extension of the original 2D point tracking model TAPIR, designed to output 3D point trajectories. It initializes keypoints in each frame, refines their 2D tracks using temporal context from adjacent frames, and subsequently lifts them to 3D space using depth estimation models such as ZoeDepth [37].
We propose a novel approach that enhances the accuracy of 3D point tracking by integrating it with the practical requirements of embedding visualizations into real-world video content. Our system enables users to anchor visualizations to fixed locations within the scene or dynamically attach them to moving objects, regardless of camera motion or scene complexity. This flexibility allows for robust visualization embedding across a wide range of video scenarios.
@@ 3 System Design and Overview
The ChartBlender system was motivated by our work experience on a data video design in collaboration with a prominent Chinese media organization. In this project, we were responsible for designing visualization charts and embedding the charts into a video. To support this task, we surveyed existing tools and systems for creating chart-embedded video. One notable system was VisTellAR [9], which enables end-users to record short videos and embed data visualizations using augmented reality (AR) techniques. VisTellAR tracks static elements in the video and anchors charts accordingly, offering a lightweight and mobile-friendly approach.
@@
Although VisTellAR is not open-source, we introduced it to our collaborators as a representative example of chart-embedded video tools. They noted, however, that end-users seldom embed charts into videos directly. In contrast, professional data journalists (our target users) require such capabilities but demand far greater precision and control. They highlighted three critical requirements: (1) accurate video motion tracking that can handle both static and dynamic objects; (2) compatibility with diverse video sources, including those lacking camera metadata, as is common in real-world media production; and (3) support for multiple dynamic charts within a single video, with fine-grained control over timing and placement to maintain contextual relevance and visual coherence.
We observed a significant gap in tools that enable seamless and efficient embedding of visualizations into real-world videos. In current practice, designers often resort to manually inserting pre-rendered charts frame by frame using software like Adobe After Effects. This process so labor-intensive that a ten-second clip can require an entire day of work according to our own experience (see the example video  [^1] http://edu.people.com.cn/n1/2023/0915/c1006-40078200.html (2’44”, 3’25”) that we produced in this way). The majority of this effort goes into repetitive adjustments of position, scale, and orientation, rather than creative or analytical tasks. This inefficiency underscores the need for a dedicated authoring tool that can automate chart-video synchronization. Through close discussions with our collaborators, we distilled a set of core design requirements that guided the development of ChartBlender.

### v2

@@ Abstract
Embedded data visualizations have emerged as a powerful narrative medium for conveying complex information within video footage. However, creating such content remains labor-intensive, as existing workflows rely on manual frame-by-frame adjustments to ensure spatial and temporal consistency. To address these challenges, we present ChartBlender, an interactive authoring system designed to streamline the creation, embedding, and automatic synchronization of data visualizations within video scenes. We develop a tracking pipeline that supports both object and camera tracking, ensuring robust alignment of visualizations with dynamic video content. To maintain visual clarity and aesthetic coherence, we also explore the design space of video-suited visualizations and develop a library of customizable templates optimized for video embedding. We evaluated ChartBlender through two controlled experiments and expert interviews with five domain experts. Results show that our system enables accurate synchronization and accelerates the production of data-driven videos.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@ 1 Introduction
Embedding data visualizations into videos has emerged as a powerful way to communicate complex information [idvxNVSurvey, amini2015understanding, shi2021autoclips, 10891192, 7539328]. This approach is now widely adopted in domains such as sports [iball], traffic [hsieh2016traffic], and journalism [covid19_youtube_2021]. By integrating charts into footage, it can provide timely and contextual insights while allowing viewers to stay focused on the underlying footage [iball]. Real-world examples of this visual medium can be seen in a video from The Economist [^1] https://www.youtube.com/watch?v=pFaGbtDJFwA&list=WL&index=8(4’35’’--5’20’’) (Fig. 1(a)). In this segment, a line chart is anchored to a wall and remains synchronized with camera motion to illustrate behavioral changes during and after the pandemic.
Figure 1: Real-world examples of data visualizations embedded in videos. (a) A frame from The Economist, where a line chart illustrates behavioral changes during and after the pandemic. (b) A frame from a national media organization in China, where a bar chart shows year-by-year growth in participant numbers.
Despite growing interest, producing videos with embedded visualizations remains a challenging task. Our collaboration with a national Chinese media organization  [^2] http://edu.people.com.cn/n1/2023/0915/c1006-40078200.html(2’44’’,3’25’’) (Fig. 1 (b)) revealed critical workflow bottlenecks. During the collaboration, we were tasked with embedding a 3D bar chart into an aerial landscape shot to visualize participant growth. The chart needed to remain synchronized with camera motion. However, to the best of our knowledge, no dedicated authoring tool provides motion-aware synchronization tailored to embedded data visualizations. In practice, we relied on general-purpose video-edit software and laborious manual keyframing of the chart’s pose and trajectory. Even minor revisions often triggered substantial rework, and producing a 10-second clip could take nearly a full workday. These bottlenecks directly informed our design requirements and system goals, highlighting an urgent need for an integrated, motion-aware authoring system for professional data-driven video production.
Within the visualization community, interest in video-embedded visualizations is also increasing. However, existing tools do not meet the requirements of motion-aware visualization authoring in real-world production. For example, sports annotation systems [7539370, iball, swimflow] demonstrate the value of integrating tracked graphics into video narratives, yet they often rely on domain-specific pipelines, fixed visual encodings, or approximate 2D overlays rather than full scene-consistent integration. Meanwhile, some AR/VR-based authoring tools [8611113, VisTellAR] enable end users to quickly place charts onto video, emphasizing ease of interaction and shareability; but they offer limited visual and stylistic control, make precise alignment and post-capture refinement difficult, and scale poorly to multi-chart compositions. Commercial video editing platforms (e.g., After Effects [adobe_ae], Blender [blender], DaVinci Resolve [davinci]) provide motion tracking and compositing capabilities, but visualization embedding typically requires labor-intensive workflows, substantial expertise, and disjoint processes that separate data design from in-context visual authoring. Overall, while motion tracking is widely available, existing tools rarely couple it tightly with visualization authoring, limiting dynamic, context-aware chart placement in complex scenes with editable visual encodings.
@@
Therefore, there is a pressing need for an interactive authoring system that supports flexible and efficient embedding of data visualizations into complex video scenes while maintaining accurate motion synchronization throughout the video. However, to design such an effective system, three challenges need to be tackled: (1) robustly tracking motion by accounting for both object movement and camera pose variation to ensure consistent synchronization between the visualization and the scene; (2) supporting generalized chart designs that remain legible and informative under dynamic video conditions; and (3) enabling simple, integrated interactions for user-driven authoring and iterative refinement. Addressing these challenges is essential for precise and seamless visualization embedding in real-world videos.
@@
To address these challenges, we present ChartBlender, an authoring system that embeds and synchronizes data visualizations with video content. Users specify a tracking point and choose a tracking mode, and the system automatically generates the motion trajectory. At any time, users can adjust both the chart’s visual style and its 3D position in the scene. In this way, the system supports the human-machine collaboration to generate data-driven videos rapidly while ensuring the visualization appears visually coherent and seamlessly embedded into the video. The major contributions of the paper are as follows.
@@
  • System. We introduce ChartBlender, a novel authoring system that enables the authoring and automatic synchronization of data visualizations with video content. It features an interactive interface that supports the end-to-end creation of visualization-embedded videos through an integrated authoring workflow. We also developed an algorithm pipeline to support robust visualization embedding for the system.
@@
  • Evaluation. We demonstrate the utility of ChartBlender through a set of example videos created from real-world data, two controlled experiments that validate the tracking pipeline, and a series of interviews with five expert users.
@@ 2 Related Work
In this section, we review the recent studies that are most relevant to our work, including embedding visualizations in video, video editing software, and point tracking.
2.1 Embedding Visualizations in Video
Embedding visualization, as a form of embedded data representation, places data-encoded graphics in close spatial correspondence with real-world referents (e.g., objects or locations), enabling viewers to interpret data in context rather than consulting a separate chart view [7539328]. When the target medium is video, this coupling should be preserved under continuous camera and scene motion, making the problem a form of visualization in motion[yao2020]. In this setting, relative motion and limited viewing time can affect how accurately and efficiently viewers read values, while professional production further requires stable spatiotemporal alignment and reliable legibility under changing viewpoints, cluttered backgrounds, and occlusions.
Early efforts such as SmartShot [smartshot] examined static embedding by optimizing chart placement to reduce occlusion of salient content and to balance aesthetic integration with readability. Beyond appearance, however, realistic embedding in dynamic 3D footage requires viewpoint-stable registration. As a result, perspective consistency, correct depth ordering, and occlusion handling become essential. To preserve visual coherence, embedded charts should follow camera trajectories and object dynamics; accordingly, a range of techniques and systems have been proposed [7539370, 10.1145/3411764.3445337, 10.1145/3411764.3445344, 10.1109/TVCG.2021.3114774]. For example, SwimFlow [swimflow] embeds animated visualizations in swimming videos by annotating key positions and interpolating trajectories, which works well for predictable motion but is less general for arbitrary scenes. iBall [iball] overlays dynamic cues in basketball videos and adapts displays based on gaze; however, under camera motion the graphics are only approximately anchored and do not provide perspective-consistent 3D registration. Overall, many existing systems rely primarily on 2D overlays and tracking, offer limited depth-aware occlusion reasoning, and are tailored to specific domains.
More recent work has explored general-purpose authoring systems that synchronize charts with arbitrary videos. With advances in AR/VR, it has become easier to create motion-following overlays. Chen et al. [8611113] introduced MARVisT, a mobile authoring tool that lets non-experts bind data to virtual glyphs and real objects, and systems such as VisTellAR [VisTellAR] enable casual users to overlay visualizations on videos quickly. However, these end-user tools typically trade expressive control and production robustness for accessibility: they provide limited chart grammars and styling controls, make precise alignment and post-capture refinement difficult, and scale poorly to multi-chart compositions.
In this work, we introduce ChartBlender as an interactive authoring system for embedding and synchronizing charts in video. ChartBlender supports 3D object tracking and perspective-consistent registration to anchor charts stably in dynamic scenes. Unlike prior systems that rely on fixed overlays or narrow workflows, ChartBlender enables flexible, in-context authoring: designers can edit chart types, styles, and data mappings directly within the video while maintaining precise control over placement and timing. This integration streamlines the creation of motion-aware embedded visualizations for professional production.
@@ 2.2 Video Editing Software
Nowadays, modern video editing and VFX tools provide motion tracking capabilities that allow designers to attach graphics or objects to dynamic scenes. After Effects [adobe_ae], for example, includes point tracking, spline-based tracking, and 3D camera tracking, enabling powerful compositing; in practice, however, achieving stable, perspective-consistent overlays, especially in the presence of fast motion or occlusions, often requires substantial manual refinement. Blender [blender] provides camera tracking and 3D scene integration within a full 3D authoring environment, but producing polished results typically requires expertise and nontrivial setup. Final Cut Pro [final_cut] offers an accessible, machine-learning-based object tracker for effects such as blurring or target-following, which is efficient for lightweight editing but provides less fine-grained control. DaVinci Resolve’s Fusion module [davinci] supports planar tracking and 3D camera tracking in a node-based workflow, yet its learning curve and context switching between editing and compositing can hinder iterative authoring. Foundry Nuke [nuke] is an industry-standard compositor with advanced planar tracking and 3D camera solving; it offers high precision but is costly and complex, and is therefore most commonly used in professional studio pipelines.
Across these tools, automatic tracking alone often falls short for chart-embedded video authoring because (1) tracking is not visualization-specific and frequently demands frame-by-frame refinement for chart stability, (2) professional workflows impose steep learning curves and high authoring overhead, and (3) chart design and tracking are decoupled, making in-context preview and iterative refinement of visualization appearance cumbersome. By contrast, ChartBlender offers a lightweight and learnable workflow with interactive controls that allow authors to adjust chart styling and placement while immediately previewing the results in context.
@@ 2.3 Point Tracking
Tracking arbitrary points in videos has witnessed substantial progress in recent years [zhou2020tracking, gauglitz2011evaluation, 10.1007/978-3-030-58548-8_28], with methods such as TAP-Net [hu2020tap] demonstrating the ability to reliably follow both static and dynamic points throughout a video sequence. These approaches typically leverage appearance-based cues and temporal consistency to maintain accurate correspondences, even under significant motion and occlusion. However, focusing solely on 2D tracking is often insufficient in practical applications. While a dynamic point in the scene can be tracked frame-by-frame, certain use cases, such as visual overlays or interactive annotations, require this point to remain anchored relative to the static scene background, rather than to moving objects or transient visual cues. To achieve this, it becomes necessary to disentangle camera motion from scene dynamics, which calls for robust camera pose estimation.
Fortunately, a variety of mature techniques now exist for estimating camera pose from monocular or multi-view video streams [xu2024critical, guan2024survey], ranging from classical SLAM-based pipelines [campos2021orb, vidal2018ultimate, 7219438] to deep learning methods[shavit2019introduction, schonberger2016structure, Hidalgo-Carrio_2022_CVPR, zhou2017unsupervised, en2018rpnet]. The integration of 2D point tracking with camera pose estimation enables a more unified paradigm: tracking arbitrary points in 3D space. This formulation has recently attracted increasing attention [koppula2024tapvid, yu2023videodoodles, xiao2024spatialtracker], as researchers extend traditional 2D tracking frameworks into the 3D domain. For instance, SpatialTracker[xiao2024spatialtracker] is the first work specifically designed for the task of 3D point tracking. It estimates the 3D trajectory of arbitrary pixels by combining image features with monocular depth estimation, and propagates information through a graph neural network (GNN) to model long-range spatiotemporal motion. Similarly, TAPIR-3D[koppula2024tapvid] is an experimental extension of the original 2D point tracking model TAPIR, designed to output 3D point trajectories. It initializes keypoints in each frame, refines their 2D tracks using temporal context from adjacent frames, and subsequently lifts them to 3D space using depth estimation models such as ZoeDepth [bhat2023zoedepth].
We propose a pipeline that enhances the accuracy of 3D point tracking by integrating it with the practical requirements of embedding visualizations into real-world video content. Our system enables users to anchor visualizations to fixed locations within the scene or dynamically attach them to moving objects, regardless of camera motion or scene complexity. This flexibility allows for robust visualization embedding across a wide range of video scenarios.
@@ 3 System Design and Overview
In this section, we present the design of ChartBlender. We first introduce our collaboration with a national Chinese media organization. Drawing on their feedback and our observations, we distill four design requirements (R1–R4) and then provide an overview of the system.
3.1 Collaboration with Professional Data Journalists
The development of ChartBlender was motivated by a collaboration with a national Chinese media organization. Our research team was invited to provide technical support for a micro-documentary on the College Student Volunteer Service Western Project, which aimed to narrate volunteers’ stories and the program’s long-term impact on regional development. We assembled a project team of two professors and four graduate students with expertise in data visualization. We worked closely with five professional data journalists from the organization (average 15 years of work experience), who were highly familiar with data videos. Using raw footage and data provided by the organization, we were tasked with producing three chart-embedded video segments that required high-precision visual integration. One segment embedded a 3D bar chart into an aerial drone shot to depict year-by-year changes in volunteer participation, with the chart appearing anchored to the terrain. The other two segments overlaid key indicators that had to remain tightly aligned with moving targets in the footage.
The collaboration followed a 15-day production cycle with daily online review meetings and rapid iterative revisions. In each meeting, our team presented the latest composites together with intermediate project files, and the journalists evaluated them shot by shot. They requested targeted adjustments such as changing anchoring targets, fine-tuning entrance and exit timing, refining visual styling for on-screen legibility, and repositioning overlays. Between meetings, we implemented these requests and circulated updated renders for asynchronous feedback, which was then consolidated and addressed in subsequent reviews. This recurring loop of authoring, critique, and revision not only ensured visual precision but also exposed where existing workflows break down, what corrections experts repeatedly perform, and which controls they consider essential for professional-grade results.
@@ 3.1 Collaboration with Professional Data Journalists
To support the task, we surveyed prior systems for chart-embedded video authoring. One notable system is VisTellAR [VisTellAR], which enables end users to capture short videos and embed data visualizations through augmented reality (AR). VisTellAR tracks static elements in the scene and anchors charts accordingly, offering a lightweight, mobile-friendly workflow. Although VisTellAR is not open source, we introduced it to our collaborators as a concrete reference for discussing chart-in-video authoring. They noted, however, that unlike end-user workflows, professional data journalists (our target users) typically embed charts during post-production and demand substantially higher precision and finer-grained control. Across production iterations and review discussions, three needs repeatedly surfaced: (1) accurate motion tracking that supports both static structures and dynamic objects; (2) compatibility with diverse footage, including videos without camera metadata, which is common in practice; and (3) support for multiple dynamic charts within a single video, with fine-grained control over timing and placement to maintain visual coherence.
In the absence of dedicated tools, we followed a typical professional workflow to produce the segments. We first created charts in 3D software, exported them as 3D assets, imported the assets into Adobe After Effects, applied automatic motion tracking to obtain an initial alignment, and then refined the results by adjusting position and orientation frame by frame before final compositing. During this process, we observed several bottlenecks that hinder seamless and efficient embedding. Although professional compositing software provides tracking functions, tracking results can be unstable in challenging shots, requiring repeated adjustments of position, scale, and orientation to maintain spatiotemporal consistency. Moreover, chart authoring and motion tracking are decoupled, preventing designers from previewing and refining charts in context while editing. In our experience, producing a 10-second clip with this workflow can take several hours, even a full workday, with much of the effort devoted to repetitive refinements rather than higher-level design work. These observations motivate a dedicated authoring tool that tightly integrates chart design with robust tracking to automate chart–video synchronization while preserving the precise control required in professional production.
3.2 Design Requirements
Through discussions with our collaborators, we distilled four design requirements (R1-R4) that guided the development of ChartBlender, which shaped both the system’s architecture and interface.